OpenStudy (anonymous):

what does 5x-1=2x+5 equal? i tried working it out and i keep getting a decimal.

jimthompson5910 (jim_thompson5910):

What did you get for the answer?

OpenStudy (anonymous):

1.3

jimthompson5910 (jim_thompson5910):

5x-1=2x+5 5x=2x+5+1 5x-2x=5+1 Did you get this as part of your work?

OpenStudy (anonymous):

the first one, 5x-1=2x+5 and everytime i work it out i get a decimal and it keeps on saying its wrong.

jimthompson5910 (jim_thompson5910):

No I meant the steps I listed out above, did you follow those same steps?

OpenStudy (anonymous):

no i did it differently

jimthompson5910 (jim_thompson5910):

can I see how you did it?

OpenStudy (anonymous):

yeah i did it two ways, just to see if i got different answers and i did and they both were a decimal and its not supposed to be a decimal. heres one. 5x-1=2x+5 3x-1=5 then i subtracted the 1 from both sides (1 and 5) and i got 4 and then i divided 4 and 3 and got 1.3

jimthompson5910 (jim_thompson5910):

ah i see where you went wrong

jimthompson5910 (jim_thompson5910):

you need to add one to both sides to get 3x-1=5 3x-1+1=5+1 3x = 6

OpenStudy (anonymous):

ahhhh i got yaaa thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome :)